### v1

# Grammatical Error Correction for Low-Resource Languages: The Case of Zarma

Abstract
Grammatical error correction (GEC) is important for improving written materials for low-resource languages like Zarma—spoken by over 5 million people in West Africa. Yet it remains a challenging problem. This study compares rule-based methods, machine translation (MT) models, and large language models (LLMs) for GEC in Zarma. We evaluate each approach’s effectiveness on our manually-built dataset of over 250,000 examples using synthetic and human-annotated data. Our experiments show that the MT-based approach using the M2M100 model outperforms others, achieving a detection rate of 95.82% and a suggestion accuracy of 78.90% in automatic evaluations, and scoring 3.0 out of 5.0 in logical/grammar error correction during MEs by native speakers. The rule-based method achieved perfect detection (100%) and high suggestion accuracy (96.27%) for spelling corrections but struggled with context-level errors. LLMs like MT5-small showed moderate performance with a detection rate of 90.62% and a suggestion accuracy of 57.15%. Our work highlights the potential of MT models to enhance GEC in low-resource languages, paving the way for more inclusive NLP tools.
Grammatical Error Correction for Low-Resource Languages: The Case of Zarma
Mamadou K. KEITA1, Christopher Homan3, Sofianou Abdoulaye Hamani1, Adwoa Bremang2, Marcos Zampieri4, Habibatou Abdoulaye Alfari2, Elysabhete Amadou Ibrahim2, Dennis Owusu2 1NILab, 2Ashesi University, 3Rochester Institute of Technology, 4George Mason University
1 Introduction
Grammatical error correction (GEC) is an essential task in natural language processing (NLP) that aims to improve the quality and readability of texts by correcting grammatical errors. GEC are important for enhancing written materials, significantly impacting educational outcomes, professional opportunities, and access to information. This is particularly relevant in under-resourced settings where limited access to academic resources and formal training can amplify language proficiency disparities. Furthermore, communities in these settings often rely on local languages for communication and knowledge transmission. While GEC tools are well-developed for high-resource languages—especially English—where extensive annotated datasets and standardized writing systems are available (Napoles et al., 2019), GEC presents significant challenges for low-resource languages.
Zarma, a Nilo-Saharan language spoken by over 5 million people across Niger and neighboring countries (Lewis et al., 2016), exemplifies the difficulties faced by many low-resource languages. The lack of standardized orthography and the limited availability of annotated data make it challenging to develop practical GEC tools. These challenges are not unique to Zarma but are shared by a broad class of low-resource languages, including many African and indigenous languages worldwide. Addressing these challenges requires innovative approaches that can work with minimal data and accommodate non-standardized texts.
The emergence of large language models (LLMs) gave notice to a new era in NLP, empowering machine learning models to learn and generate human-like text across many languages (Devlin et al., 2018; Brown et al., 2020). LLMs exhibit remarkable zero-shot and few-shot learning capabilities (Wan et al., 2023), which are potentially beneficial for low-resource languages. These capabilities allow LLMs to perform effectively with minimal data, making them valuable for languages with limited annotated datasets. However, the use of LLMs is also challenging in low-resource settings. They are primarily trained on data from high-resource languages, which may limit their performance when applied to low-resource languages. The lack of representative training data can lead to errors and biases, reducing their effectiveness. Additionally, the significant computational resources required to fine-tune and deploy LLMs can be a barrier in resource-constrained environments.
This study investigates the potential of LLMs and traditional models to improve GEC for Zarma by comparing conventional rule-based methods, machine translation (MT)-based models, and the novel application of LLMs for Zarma GEC. To explore the performance of these models on other low-resource languages, we replicate our Zarma GEC experiments with Bambara, a West African language spoken in Mali. Through a case study, we highlight the strengths and weaknesses of each method, aiming to bridge the linguistic gap in NLP for low-resource languages like Zarma and Bambara.
Our research addresses the following questions:
RQ1:
Do state-of-the-art LLM models outperform conventional rule-based and MT-based models on GEC for Zarma texts?
RQ2:
What are the specific strengths and limitations of each GEC approach in low-resource settings, considering the lack of standardized orthography and limited annotated data?
The main contributions of this paper are as follows:
  • A comprehensive evaluation of three distinct GEC approaches for Zarma (rule-based, MT-based, and LLM-based). Our findings show that the MT-based approach delivered the highest accuracy with a detection rate of 95.82%, a suggestion accuracy of 78.90%, and acceptable performance in zero-shot scenarios.
  • Reproduction of the experiments with Bambara—a manding family language—to confirm replicability and broaden the study’s scope beyond Zarma.
  • Development and planned public release (upon acceptance) of models fine-tuned for GEC of the tested languages.
2 Related Work
Cissé and Sadat (2023) present an approach for spellchecking Wolof, a language primarily spoken in Senegal. Their algorithm combines a dictionary lookup with an edit distance metric Levenshtein distance algorithm (Levenshtein, 1966) to identify and correct spelling errors.
Vydrin and collaborators developed Daba, a software package for grammar and spellchecking in Manding languages (Vydrin, 2014). Their work employs a rule-based system focusing on morphological analysis, addressing the agglutinative nature of Manding languages.
Researchers have explored the use of LLMs for language-specific tasks, including GEC, demonstrating their adaptability beyond high-resource languages. For instance, a study by Palma Gomez et al. (2023) showed that the MT5 model (Xue et al., 2020), a multilingual transformer model pre-trained on a massive dataset of text and code, could be effectively fine-tuned for GEC in Ukrainian. Song et al. (2023) present an innovative application of LLMs—specifically GPT-4 (OpenAI, 2023)—for generating explanations for grammatical errors. Another promising approach to GEC leverages pre-trained multilingual MT models. The study by (Luhtaru et al., 2024) introduced the "No Error Left Behind" approach, which uses models like MT5 and NLLB (Team et al., 2022). The fine-tuning process involves adapting the pre-trained MT model to treat error correction as a "translation" task, where the source language is the incorrect sentence and the target language is the corrected sentence. However, the research also identified a fundamental challenge: the trade-off between precision and recall when training with synthetic data. This finding shows the need for further research to optimize these models for low-resource GEC tasks, potentially through improved data augmentation techniques or developing specialized architectures for error correction.
3 Proposed Methods
3.1 GEC with LLMs
LLMs have significantly advanced NLP, exhibiting capabilities in multitasking, few-shot and multilingual learning. These models, extensively pre-trained on diverse datasets, demonstrate a remarkable ability to grasp nuanced aspects of language, reasoning, and context (Brown et al., 2020; Raffel et al., 2020). This section outlines our methodology for leveraging LLMs to develop a GEC tool for Zarma. The proposed GEC tool is designed to function independently of predefined grammar rules or lexicons, using the models’ few-shot learning capabilities for enhanced efficiency in low-resource scenarios.
3.1.1 Implementation
We employed two distinct approaches for LLM-based GEC:
Instruction and Error Explanation Fine-tuning
This method involves embedding training data within a contextual sentence structure, enhancing the model’s reasoning abilities, and facilitating effective learning from the examples. We use the following instruction prompt:
Prompt: "Zarma sentence: [Incorrect Sentence],Correct the zarma sentence: [Correct Sentence]Error Causes: : [Error $\begin{array}[]{l}\text{{Prompt}: "Zarma sentence: [{Incorrect Sentence}],}\\ \text{Correct the zarma sentence: [{Correct Sentence}]}\\ \text{{Error Causes: }: [{Error Cause}]."}\end{array}$
This format, particularly the error explanation component, is important for aiding the model’s comprehension of Zarma’s grammatical contexts and patterns, as demonstrated in previous research (Schick and Schütze, 2021; Wei et al., 2021).
Non-Prompt Fine-tuning Using Aligned Sentences
This approach involves fine-tuning the model directly on parallel data of incorrect and correct Zarma sentences without explicit prompts. This leverages the model’s ability to learn the implicit mapping between incorrect and correct forms.
3.2 GEC with MT Models
Exploring MT models for GEC represents a promising avenue for addressing GEC challenges in low-resource languages like Zarma. MT models, particularly those pre-trained on multilingual datasets, offer strong capabilities for processing text across diverse linguistic frameworks.
3.2.1 Implementation
We chose the M2M100 model (Fan et al., 2020) for its demonstrated ability to translate between many languages, indicating its potential to capture cross-linguistic patterns relevant to GEC. To adapt M2M100 for Zarma GEC, we fine-tuned it on a corrupted corpus designed to reflect common errors in Zarma text. This corpus was generated by applying a custom noise script, described in Section 4. The script introduces various errors, ensuring the training data effectively represents realistic challenges in real-world Zarma text. This corrupted corpus, paired with the original correct sentences, serves as the training data for M2M100, enabling the model to learn the mapping from incorrect to correct Zarma. Detailed training settings and evaluation metrics for this MT-based GEC approach can be found in Tables 7, 4, and 5.
3.3 Rule-based GEC for Zarma
As a baseline, we designed a rule-based GEC process, based on Levenshtein distance and the Bloom filter (Bloom, 1970), for Zarma (Figure 1). Additionally, we implemented a tool and API in Python. The tool—the first of its kind for Zarma, to our knowledge—is designed to cater to a wide range of users. It offers command-line and graphical user interfaces (GUI) and is much less computationally intensive than the LLM-based approach. Moreover, our results show that it provides a competitive spell correction performance compared to the LLMs- and MT-based approaches. To ensure the tool’s accessibility, we plan a public release upon acceptance.
Figure 1: Rule-Based GEC tool Workflow
The GEC process for Zarma text follows a structured sequence of steps, starting with the input text $T$, passing through a Bloom filter decision $B$ to determine if the word exists in the dictionary, leading to either a grammatical check or a correction algorithm, followed by a final dictionary lookup and correction $C$, as needed:
$C(T)=\begin{cases}T&\text{if }D(B(T))=\text{True},\\ \text{Correct}(T)&\text{if }D(B(T))=\text{False}.\end{cases}$ (1)
The detailed process includes:
Text Processing
The system begins by processing the text, where the Zarma content is segmented into words, punctuation, and spaces using regular expressions. This initial step ensures that every part of the text is ready for individual examination—a necessity given Zarma’s linguistic intricacies.
Bloom Filter
The Bloom filter—known for its space and time efficiency—performs probabilistic checks on whether a word might be in the dictionary. For a given word $w$, it uses a set of hash functions to probe various positions in a bit array. If all checked positions are flagged, $w$ might be in the dictionary:
$B(w)=\bigwedge_{i=1}^{k}\text{bit}[h_{i}(w)],$ (2)
If $B(T)=\text{False}$, the system assumes the word is not in the dictionary and immediately moves to the Correction Algorithm to suggest a correction. This process ensures efficiency when dealing with unknown words.
Dictionary Lookup
At the core of the system is the trie-based dictionary lookup. If the Bloom Filter decision is positive ($B(T)=\text{True}$), the system proceeds to a Grammatical Check to verify if the word is grammatically correct. If the word is found to be grammatically incorrect, or if $B(T)=\text{False}$ (indicating the word is not in the dictionary), the system performs a Dictionary Lookup to confirm or correct the word.
Grammar Rules
The system incorporates a set of Zarma grammar rules, a task made challenging by the lack of a standard writing format. This phase scrutinizes elements like consonant rules and vowel lengths, essential for keeping the text accurate to Zarma’s grammatical essence despite its absence of a writing standard.
Correction Algorithm
When an error is detected, either because the Bloom Filter returned false or the grammatical check identified an issue, the Correction Algorithm is activated. This component leverages the Levenshtein distance $L$ to identify the smallest number of edits—insertions, deletions, or substitutions—required to rectify the erroneous word. For illustration, given a corrupted sentence "A sindq biri," the Levenshtein algorithm’s operation can be represented as:
Corrupted Sentence: "A sindq $\begin{array}[]{c}\text{{Corrupted Sentence}: "A sindq biri"}\\ \downarrow\\ \text{{LevSuggest}}("sindq")\\ \downarrow\\ \{\text{"sind"},\text{"sinda"}\}\end{array}$ (3)
Here, "sindq" undergoes comparison with the lexicon, and "sind" and "sinda" are suggested based on their proximity in terms of minimal edit distance. In this context, $L(a,b)$ calculates the distance between the incorrect string $a$ and a potential correct string $b$, considering their lengths $i$ and $j$, with $1_{(a_{i}\neq b_{j})}$ serving as an indicator function to highlight discrepancies. This mechanism ensures the algorithm’s efficiency in providing appropriate corrections, thus enhancing the GEC tool’s overall accuracy for Zarma text.
4 Data Curation Insights
In this section, we detail our approach to gathering and preparing data for training our GEC models for Zarma. We constructed two distinct datasets—a synthetic dataset and a human-annotated “gold” dataset—which we then combined to create a comprehensive training set. The final dataset comprises over 250,000 examples, split into 80% for training, 10% for validation, and 10% for testing.
4.0.1 Synthetic Data
To address the scarcity of annotated data in Zarma, we generated a synthetic dataset by applying a custom corruption script to the Feriji Dataset (Keita et al., 2024). This script introduces realistic typographical and grammatical errors into Zarma sentences, mimicking plausible human mistakes. The corruption process is guided by a defined set of operations to ensure the errors are representative of actual language use.
Specifically, we applied a noise function N to each original sentence SZ (composed of a subject S, verb V, and complement C) to produce a corrupted sentence CSZ:
$CSZ(S,V,C)=N\circ SZ(S,V,C)$ (4)
The noise function N consists of four operations: deletions $\delta$, insertions $\mu$, substitutions $\sigma$, and transpositions $\tau$. These operations are designed to introduce errors without inadvertently creating another grammatically valid Zarma sentence.
For example, consider the original sentence SZ = A go koy fuo” (He is going home”). Applying a substitution operation $\delta$ that changes ’go’ to ’ga’ yields the corrupted sentence CSZ = ’A ga koy fuo’. While ’ga’ is a valid Zarma word, in this context, it alters the meaning to ’He will go home,’ introducing a grammatical error.
$\textbf{SZ}="A\,go\,koy\,fuo"$ (5)
$\textbf{N}(\textbf{SZ})=\textbf{SZ}\xrightarrow{\sigma("go"\to"ga")}"A\,ga\,koy\,fuo"$
$\textbf{CSZ}="A\,ga\,koy\,fuo"$
[…]
Furthermore, to enrich the dataset with diverse linguistic nuances, the script generates four corrupted variants for each correct sentence in the dataset. A snapshot of the synthetic data is provided in Table 1.
Correct Sentence Corrupted Sentences
Sintina gaa Irikoy na beena da ganda taka. Sintina gaa Irikog na beena da ganda taka.
Sintina gaa Irikoy na been da ganda taka.
Sintina aga Irikoy na beena da ganda taka.
Sintina gaa Irikoy na beena ea ganda taka.
Table 1: Snapshot of The Synthetic Data
4.1 Human-Annotated Data
The second dataset, referred to as the ‘gold data,’ was curated by human annotators proficient in Zarma. They introduced grammatical and logical errors (LG)—more details about this errors type can be found in Section B.2—into Zarma sentences and provided justifications for each alteration. This dataset exposes the models to a broader array of error types—particularly LG errors—and their corresponding corrections, enhancing the models’ capacity to generalize and accurately correct unseen texts in Zarma, especially in zero-shot or few-shot learning scenarios. An example from the gold data is shown in Table 2
Incorrect Sentence Correction Error Explanation
Souba, Ay koy Niamey Souba, Ay ga koy Niamey "Souba" means tomorrow, and therefore the tense must be in the future using the future tense marker "ga" after the subject "Ay."
Table 2: Snapshot of the Gold Data
5 Experiment
We conducted a series of experiments to evaluate the effectiveness of different GEC methods for Zarma. We selected three models for training based on their demonstrated proficiency in multilingual tasks and aptitude for few-shot learning: Gemma (Team et al., 2024), MT5-small, and M2M100. The training was conducted on Google Cloud, using a Deep Learning Virtual Machine equipped with NVIDIA GPUs T4 x2. Due to computational resource constraints, we applied QLoRA quantization (Dettmers et al., 2023) to the Gemma model to reduce memory requirements. For MT5-small and M2M100, we used smaller variants of the models to ensure feasible training times. For the LLM-based methods (Gemma and MT5-small), we structured the combined dataset as described in Section 4, incorporating both incorrect sentences and their corrections along with error explanations. For the MT-based method (M2M100), we adopted a MT task-specific format, using aligned pairs of incorrect and correct Zarma sentences without explicit error explanations. The detailed training settings for each model are presented in Table 7.
6 Results & Comparative Analysis
To assess the effectiveness of each GEC method for the types of errors addressed in this paper (as detailed in Section B), we conducted a comprehensive evaluation consisting of two parts: an automatic evaluation (AE) and a manual evaluation (ME) by native Zarma speakers. We included the ME because AE tools are primarily developed and tested on high-resource languages. Therefore, their scores may not reliably reflect performance in low-resource settings like Zarma. The ME allows us to obtain assurances from language speakers regarding the actual effectiveness of the methods.
6.1 Automatic Evaluation Results
In the AE, we employed two sets of metrics to evaluate our techniques. The first set consisted of common machine learning metrics, including the GLEU score (Wang et al., 2018) and the M2 score (Dahlmeier and Ng, 2012). However, these metrics are not applicable to the rule-based method due to its deterministic nature and the way these metrics are designed, which favor probabilistic models that produce varying outputs. The second set focused on spelling correction metrics—measuring the detection rate and suggestion accuracy. The results are presented in Table 3.
Methods AE Evaluation
Common ML Metrics Spelling Correction
GLEU Detection Correction
Rule-based - - 100% 96.27%
Gemma 2b 0.4572 0.528 76.19% 43.28%
MT5-small 0.6203 0.731 90.62% 57.15%
M2M100 0.7896 0.914 95.82% 78.90%
Table 3: AE metrics scores
6.1.1 Common Machine Learning Metrics
The GLEU and M2 scores provide an overall assessment of the models’ performance in correcting grammatical errors by comparing the corrected sentences to reference sentences. The higher scores achieved by the M2M100 model (GLEU: 0.7896, M2: 0.914) indicate its higher performance among the tested models. The LLM-based models, Gemma 2b and MT5-small, obtained moderate scores, reflecting their ability to correct some errors but not as effectively as the M2M100 model.
6.1.2 Spelling Correction Metrics
For the spelling correction evaluation, we used a set of 3,539 sentences from the Feriji Dataset, introducing typographical errors using our corruption script (as described in Section 4). We measured two key metrics:
  • Detection Rate: The percentage of errors that the method successfully identified and attempted to correct.
  • Suggestion Accuracy: The percentage of corrections suggested by the method that were accurate.
As shown in Table 3, the rule-based method achieved a detection rate of 100% and a suggestion accuracy of 96.27%. This impressive performance is primarily due to its extensive dictionary and predefined rules, which allow it to effectively flag and correct errors covered within its scope. The M2M100 model also performed well, with a detection rate of 95.82% and a suggestion accuracy of 78.90%. This suggests that the MT-based approach is capable of effectively identifying and correcting spelling errors, even in a low-resource language setting. The LLM-based models, Gemma 2b and MT5-small, showed lower detection rates and suggestion accuracies. Gemma 2b detected 76.19% of errors with a suggestion accuracy of 43.28%, while MT5-small detected 90.62% of errors with a suggestion accuracy of 57.15%. These results indicate that while LLMs have some capacity to correct spelling errors, they are less effective than the MT-based approach and the rule-based method in this specific task.
6.2 Manual Evaluation Results
Given the limitations of AE metrics in low-resource settings, we conducted the ME involving five native Zarma speakers. The evaluators assessed how well each method corrected LG errors, using a scale from 1 to 5, where 1 indicates a poor correction and 5 indicates an excellent correction. The evaluation was performed on a set of 300 sentences from the Feriji Dataset, which were manually corrupted by our annotators to include LG errors. Detailed information about the evaluation protocol and the recruitment of evaluators can be found in Section C. The results are presented in Table 4.
Methods ME avg (1-5)
Logical Errors Correction Sentence Improvement
Rule-based 0.4 0
Gemma 2b 1 0
MT5-small 1.7 1
M2M100 3 2.5
Table 4: Manual Evaluation
The rule-based method struggled with LG errors, achieving low scores (0.4 for logical error correction and 0.0 for sentence improvement). This outcome is expected, as the rule-based approach relies on predefined rules and lacks the ability to understand the context or semantics of sentences. The M2M100 model outperformed the other methods, achieving higher scores in both logical error correction (3.0) and sentence improvement (2.5). This demonstrates that the MT-based approach is more effective at handling LG errors and improving the overall quality of the sentences. The LLM-based models achieved moderate scores, with MT5-small performing better than Gemma 2b. MT5-small received scores of 1.7 for logical error correction and 1.0 for sentence improvement, suggesting some capability in handling LG errors, but not as effectively as the M2M100 model. We observed that the models faced varying levels of difficulty with different types of LG errors. For instance, they were generally better at correcting verb tense errors than subject-verb agreement errors.
6.3 Zero-Shot scores
To further assess the generalization capabilities of our models, we conducted a zero-shot evaluation using a separate dataset containing sentences with words and grammatical structures not present in the training data. This evaluation is crucial for understanding how well the models can handle new or unseen data, which is a common scenario in low-resource language settings. The results of the zero-shot evaluation are presented in Table 5.
Methods AE Evaluation Manual Evaluation avg (1-5)
Common ML Metrics Spelling Correction Logical Errors Correction Sentence Improvement
GLEU Detection Correction
Rule-based - - 94.17% 81.04% 0 0
Gemma 2b 0.3926 0.481 53.93% 26.52% 0.6 0
MT5-small 0.5018 0.657 73.64% 58.31% 0.8 1
M2M100 0.7683 0.904 92.27% 78.34% 2.7 1.3
Table 5: Performance Evaluation on Zero-Shot Dataset
As expected, the rule-based method showed a decrease in performance compared to the controlled evaluation. Its detection rate dropped to 94.17% and suggestion accuracy to 81.04%. This decline is due to the method’s reliance on a predefined dictionary and rules, which may not cover new words or grammatical constructions present in the zero-shot dataset. Additionally, the ME scored it with 0 points for both logical errors correction and sentence improvement.
The M2M100 model again demonstrated superior performance among the tested models in the zero-shot setting. It achieved a GLEU score of 0.7683 and an M² score of 0.904, indicating strong overall correction capability. Its detection rate was 92.27% with a suggestion accuracy of 78.34%, showing that it can effectively identify and correct errors in unseen data. Furthermore, in ME, it received 2.7 for logical errors correction and 1.3 for sentence improvement.
The LLM-based models, Gemma 2b and MT5-small, showed reduced performance in the zero-shot evaluation. Gemma 2b’s detection rate was 53.93% with a suggestion accuracy of 26.52%, while MT5-small achieved a detection rate of 73.64% and a suggestion accuracy of 58.31%. In terms of ME, Gemma 2b scored 0.6 for logical errors correction and 0 for sentence improvement, while MT5-small achieved 0.8 and 1, respectively.
Our results strongly suggest that MT models trained on extensive multilingual datasets hold significant potential for improving GEC in low-resource languages. The ME scores further validate that the M2M100 model excels not only in automated metrics but also in human judgments, particularly in complex correction tasks. To validate the reproducibility and robustness of our methods, we conducted additional experiments with Bambara—a language from a different linguistic family. The results of this experiment are detailed in Section A.
7 Strengths and Limitations of Each Method
7.1 Rule-Based Method
Rule-based approaches demonstrated high effectiveness in addressing predictable and previously encountered error patterns. In controlled tests, these methods achieved perfect detection and high suggestion accuracy. Their strength lies in the reliance on a comprehensive set of predefined rules and a detailed lexicon. However, this reliance also presents significant limitations—inflexibility and lack of adaptability. Rule-based methods struggle to handle new or unexpected errors, especially in dynamic language use where novel words or grammatical constructions are common. This limitation is particularly pronounced in zero-shot scenarios, where the system encounters words or structures not included in its rules or dictionary. Moreover, the dependency on extensive and carefully curated linguistic resources restricts the scalability of rule-based methods, especially for low-resource languages like Zarma, where such resources are often limited or incomplete (Scannell, 2007).
7.2 LLMs Method
The LLMs in our experiments—Gemma 2b and MT5-small—demonstrated moderate performance in both controlled and zero-shot scenarios. A key strength of LLMs is their capacity to understand context and generate corrections based on broader textual cues, rather than relying only on direct matches to known errors. However, their performance is significantly dependent on the diversity and quality of the training data. A major limitation is that most pre-existing LLMs are primarily trained on data from high-resource languages—mainly Western languages. Consequently, their applicability to African languages like Zarma is hindered by a lack of representative training examples (Bender et al., 2021). This results in lower suggestion accuracy and difficulties in effectively handling the linguistic complexities of these languages. Furthermore, training and fine-tuning LLMs require substantial computational resources, creating a significant barrier in resource-constrained environments.
7.3 MT Method
The MT approach—using the M2M100 model—demonstrated exceptional performance, particularly in zero-shot scenarios, surpassing both rule-based methods and LLMs. The strength of this approach lies in the ability of MT models to leverage multilingual translation mechanisms, effectively adapting to the nuances of diverse languages through exposure to vast and varied training datasets. This characteristic makes MT models particularly suitable for GEC in low-resource languages, as they can infer corrections from patterns learned across multiple languages. This aligns with research highlighting the effectiveness of MT models for cross-lingual transfer learning in various NLP tasks (Conneau et al., 2018). However, a significant challenge in using MT models for GEC in low-resource languages is the frequent scarcity of quality parallel corpora for training. Without sufficient data, the models may generate less accurate or contextually inappropriate corrections (Tiedemann, 2020). Additionally, despite their strengths, MT models require fine-tuning and continuous updating to maintain their accuracy and relevance, as language use evolves.
8 Potential Applications
Our team visited Niamey to present our work to the local Zarma community and gather their feedback. The discussions provided valuable insights into potential applications for our GEC tool and broader language models, highlighting their potential impact on the community.
Content Creation and Educational Resources
One significant comment we received was the potential use of our model to translate coding content and general educational materials for enthusiasts and students. There is a growing interest in technology and programming within the community, but a considerable language barrier exists due to the lack of educational resources in Zarma. By translating and correcting coding tutorials, textbooks, and other educational materials into Zarma, our model can help overcome this challenge. Additionally, the GEC tool can be used to improve general educational content in Zarma, including textbooks, instructional materials, and online courses across various subjects.
Cultural Preservation and Documentation
Feedback from the community highlighted the importance of maintaining accurate written records of folklore, oral histories, and traditional knowledge. The GEC tool can support these efforts by providing a reliable means for transcribing and publishing grammatically accurate texts.
9 Conclusion and Future Work
In this study, we have demonstrated the potential of leveraging LLMs and MT models to address the critical need for effective grammatical error correction GEC tools in low-resource languages, with a focus on Zarma. The challenges inherent to developing GEC tools for Zarma—such as data scarcity and lack of standardized orthography—necessitated innovative approaches beyond those used in high-resource language contexts. To overcome these challenges, we employed a novel methodology combining three key components: (1) the development of a custom corruption script to generate synthetic training data, thereby mitigating the limited availability of annotated Zarma text; (2) the creation of a human-annotated "Gold Data" set, incorporating expert knowledge of Zarma grammar to provide a valuable benchmark for evaluating model performance on complex errors; and (3) the adaptation and fine-tuning of advanced LLMs and MT models—Gemma 2b and M2M100—for the task of Zarma GEC.
This research establishes a baseline for future work in developing GEC tools for Zarma and other low-resource languages. Future research directions include:
Hybrid Approaches
Exploring hybrid models that combine the precision of rule-based methods with the adaptability of LLMs and the robustness of MT models could yield more effective GEC tools.
Data Augmentation
Further expanding and diversifying the training dataset, possibly through community-driven data collection and annotation efforts, to enhance model performance and generalization capabilities.
Continuous Learning Mechanisms
Incorporating mechanisms for continuous learning and user feedback could enable the models to adapt to evolving language use, idiomatic expressions, and regional dialects.
Resource Optimization
Investigating methods to reduce the computational requirements of advanced models, such as model compression or distillation techniques, to make GEC tools more accessible in resource-constrained environments.
Cross-Lingual Transfer Learning
Applying the insights gained from this study to other low-resource languages, leveraging cross-lingual transfer learning to accelerate the development of GEC tools across diverse linguistic contexts.
Appendix A Further Experiment with other Languages
After obtaining promising results for Zarma using LLM and MT-based approaches, we conducted further experiments to validate the reproducibility of these methods—using the M2M100 and MT5 models. We selected the Bambara language for this experiment because it belongs to a different linguistic family, allowing us to evaluate the performance of the approaches on a language outside the Nilo-Saharan family. We used the Bayelemabaga dataset (Vydrin et al., 2022) for Bambara. The same data preparation process described in the methodology section was followed; however, we excluded any human-annotated data to focus only on spelling correction performance. The results are presented in Table 6.
Methods AE Evaluation
Common ML Metrics Spelling Correction
GLEU Detection Correction
MT5-small 0.5713 0.643 87.45% 52.91%
M2M100 0.7293 0.851 94.64% 68.18%
Table 6: Performance Metrics for Bambara
The Bambara experiment demonstrated that the MT-based approach outperformed the LLMs-based one regarding word-level correction metrics. The MT-based approach achieved a detection rate of 94.64% and a suggestion accuracy of 68.18%. In contrast, the LLMs-based approach detected 87.45% of errors and suggested corrections with 52.91% accuracy. The promising results from the Bambara experiment highlight the potential of both LLMs and MT models to improve GEC for low-resource languages significantly. However, they also emphasize the necessity for continued expanding and diversifying training datasets.
Appendix B Errors Being Addressed
In this section, we explain the types of errors GEC methods address. We categorize the errors into two main types: word-level correction (spellchecking) and context-level correction. The context-level correction is further divided into grammar errors, logical errors, and sentence improvement. Below, we define each error type and provide examples to illustrate them.
B.1 Word-Level Correction (Spellchecking)
Word-Level correction involves identifying and correcting typographical errors in individual words. These errors are usually due to misspellings, incorrect usage of characters, or typographical mistakes.
  • Example:
    • Incorrect: Sintina gaa Irikog na beena da ganda taka.
    • Correct: Sintina gaa Irikoy na beena da ganda taka.
B.2 Context-Level Correction
Context-level correction involves errors that go beyond individual words and affect the overall structure and meaning of the sentence. We categorize these errors into logical errors and sentence improvement.
B.2.1 Logical Errors
Logical errors include incorrect verb conjugations, subject-verb agreement issues, incorrect use of grammatical markers, and logical inconsistencies within the sentence. These errors affect the grammatical correctness and logical coherence of the sentence.
  • Example:
    • Incorrect: Souba, Ay koy Niamey. (The time indicator "Souba" means "tomorrow," but the verb "koy" indicates present tense.)
    • Correct: Souba, Ay ga koy Niamey. (The future tense marker "ga" matches the time indicator "Souba.")
B.2.2 Sentence Improvement
Sentence improvement involves enhancing the quality of the sentence by making it more precise, concise, or stylistically appropriate. This category addresses grammatically correct sentences that can be improved for better readability or style.
  • Example:
    • Original: I girbi honkuna i tun be.
    • Improved: I ga girbi suba.
Appendix C Evaluators Recruitment Process
To ensure a thorough and culturally appropriate manual evaluation of our GEC methods, we recruited five native Zarma speakers to serve as evaluators. The primary criteria for selection was their fluency in Zarma—their ability to read and write proficiently in the language. The evaluators were selected through community outreach in Niamey, Niger, leveraging connections with local educational institutions. All selected evaluators had completed secondary education or higher.
Prior to the evaluation, we conducted a training workshop to familiarize the evaluators with the objectives of the study, the evaluation protocols, and the scoring guidelines. During the evaluation process, each evaluator independently assessed a set of 300 sentences that had been manually corrupted to include logical and grammatical errors. They reviewed the corrections made by each GEC method and assigned scores based on accuracy and fluency, where:
  • 1 indicates a poor correction with significant errors remaining or introduced.
  • 5 indicates an excellent correction with all errors appropriately addressed and the sentence reading naturally.
To ensure the reliability of the evaluations, we measured inter-annotator agreement using Fleiss’ kappa coefficient. The resulting kappa value was 0.74, indicating substantial agreement among the evaluators. This suggests that the scoring guidelines were clear and that the evaluators applied them consistently.
All evaluators provided informed consent to participate in the study and were compensated for their time and contributions.
Appendix D Training Details
Models Parameters Training Details
QLoRA GPU Used Lr Loss
Gemma 2b 2 billion Applied NVIDIA P100 $2\times 10^{-4}$ 1.2613
MT5-small 300 million Not Applied NVIDIA T4x2 $2\times 10^{-5}$ 0.0345
M2M100 418 million Not Applied NVIDIA P100 $2\times 10^{-5}$ 0.0214
Table 7: Training Settings for the models
Figure 2: Images of the different GEC tool interfaces. The rule-based on the left and the other approaches on the right