### v2

# Grammatical Error Correction for Low-Resource Languages: The Case of Zarma

Abstract
Grammatical error correction (GEC) aims to improve quality and readability of texts through accurate correction of linguistic mistakes. Previous work has focused on high-resource languages, while low-resource languages lack robust tools. However, low-resource languages often face problems such as: non-standard orthography, limited annotated corpora, and diverse dialects, which slows down the development of GEC tools. We present a study on GEC for Zarma, spoken by over five million in West Africa. We compare three approaches: rule-based methods, machine translation (MT) models, and large language models (LLMs). We evaluated them using a dataset of more than 250,000 examples, including synthetic and human-annotated data. Our results showed that the MT-based approach using M2M100 outperforms others, with a detection rate of 95. 82% and a suggestion accuracy of 78. 90% in automatic evaluations (AE) and an average score of 3.0 out of 5.0 in manual evaluation (ME) from native speakers for grammar and logical corrections. The rule-based method was effective for spelling errors but failed on complex context-level errors. LLMs—MT5-small—showed moderate performance. Our work supports use of MT models to enhance GEC in low-resource settings, and we validated these results with Bambara, another West African language.
Grammatical Error Correction for Low-Resource Languages: The Case of Zarma
Mamadou K. KEITA1, Christopher Homan3, Marcos Zampieri4, Adwoa Bremang2, Habibatou Abdoulaye Alfari2, Elysabhete Amadou Ibrahim2, Dennis Owusu2 1PAAIR Institute, 2Ashesi University, 3Rochester Institute of Technology, 4George Mason University
1 Introduction
Grammatical error correction (GEC) is important to improving text quality in different scenarios. High-resource languages have well-developed GEC tools (Napoles et al., 2019), but low-resource languages have few resources. Zarma, a Nilo-Saharan language spoken widely in West Africa (Lewis et al., 2016), faces these challenges.
Zarma texts lack standardized orthography and large annotated datasets. Developers need approaches that handle minimal resources and non-standard writing. Large language models (LLMs) hold promise—as they can handle various languages and tasks (Brown et al., 2020)—but performance on low-resource languages can be limited. They rely heavily on data, which is scarce for Zarma.
We investigate three GEC approaches for Zarma: a rule-based method, a machine translation (MT)-based model, and LLMs. We also test with Bambara, another West African language, to confirm generalization. This paper addresses two key questions:
RQ1:
Do LLMs outperform traditional rule-based and MT models in Zarma GEC?
RQ2:
What are the strengths and limitations of each approach, given the non-standard orthography and limited data?
RQ3:
Does the best-performing approach replicate for another low-resource language—Bambara?
The main contributions of this paper are as follows:
  • A comparison of three GEC methods for Zarma. The MT-based approach—M2M100—provided the best results, with a detection rate of 95. 82% and a suggestion accuracy of 78. 90% in automatic evaluations (AE) and an average score of 3.0 out of 5.0 in manual evaluation (ME).
  • A dataset of more than 250,000 synthetic and human-annotated Zarma examples.
  • Experiments with Bambara to confirm replicability.
  • Planned release of trained models to promote future research.
2 Related Work
Cissé and Sadat (2023) present an approach for spellchecking Wolof, a language primarily spoken in Senegal. Their algorithm combines a dictionary lookup with an edit distance metric Levenshtein distance algorithm (Levenshtein, 1966) to identify and correct spelling errors.
Vydrin and collaborators developed Daba, a software package for grammar and spellchecking in Manding languages (Vydrin, 2014). Their work employs a rule-based system focusing on morphological analysis, addressing the agglutinative nature of Manding languages.
Researchers have explored the use of LLMs for language-specific tasks, including GEC, demonstrating their adaptability beyond high-resource languages. For instance, a study by Palma Gomez et al. (2023) showed that the MT5 model (Xue et al., 2020), a multilingual transformer model pre-trained on a massive dataset of text and code, could be effectively fine-tuned for GEC in Ukrainian. Song et al. (2023) present an innovative application of LLMs—specifically GPT-4 (OpenAI, 2023)—for generating explanations for grammatical errors. Another promising approach to GEC leverages pre-trained multilingual MT models. The study by (Luhtaru et al., 2024) introduced the "No Error Left Behind" approach, which uses models like MT5 and NLLB (Team et al., 2022). The fine-tuning process involves adapting the pre-trained MT model to treat error correction as a "translation" task, where the source language is the incorrect sentence and the target language is the corrected sentence. The complexity of balancing precision and recall on synthetic data remains an open problem.
3 Proposed Methods
3.1 GEC with LLMs
LLMs have significantly advanced NLP, showing capabilities in multitasking, single-shot, and multilingual learning. These models, extensively pre-trained on diverse datasets, demonstrate a remarkable ability to grasp nuanced aspects of language, reasoning, and context (Brown et al., 2020; Raffel et al., 2020). This section outlines our methodology for using LLMs to develop a GEC tool for Zarma. The proposed GEC tool is designed to function independently of predefined grammar rules or lexicons, using the models’ few-shot learning capabilities for enhanced efficiency in low-resource scenarios.
3.1.1 Implementation
We used two distinct approaches for LLM-based GEC:
Instruction and Error Explanation Fine-tuning
This method involves embedding training data within a contextual sentence structure. We embed examples in a prompt format:
Prompt: "Zarma sentence: [Incorrect Sentence],Correct the zarma sentence: [Correct Sentence]Error Causes: : [Error $\begin{array}[]{l}\text{{Prompt}: "Zarma sentence: [{Incorrect Sentence}],}\\ \text{Correct the zarma sentence: [{Correct Sentence}]}\\ \text{{Error Causes: }: [{Error Cause}]."}\end{array}$
This format helps the model reason about corrections and reflect on why errors occur, as demonstrated in previous research (Schick and Schütze, 2021; Wei et al., 2021).
Non-Prompt Fine-tuning Using Aligned Sentences
We also fine-tuned LLMs on parallel incorrect and correct sentences. This leverages the model’s ability to align corrupted text with corrected forms.
3.2 GEC with MT Models
Using MT models for GEC represents a promising avenue for addressing GEC challenges in low-resource languages such as Zarma. MT models—particularly those pre-trained on multilingual datasets—offer strong capabilities for processing text across diverse linguistic frameworks.
3.2.1 Implementation
We chose the M2M100 model (Fan et al., 2020) for its demonstrated ability to translate between many languages, indicating its potential to capture cross-linguistic patterns relevant to GEC. We adapt M2M100 for GEC by training it to translate corrupted Zarma text into corrected versions. This approach uses a synthetic corpus generated with a noise script (Section 4). The script introduces various errors, ensuring that the training data effectively represent realistic challenges in real-world Zarma text. This corrupted corpus, paired with the original correct sentences, serves as the training data for M2M100, allowing the model to learn the mapping from incorrect to correct Zarma. Detailed training settings and evaluation metrics for this MT-based GEC approach can be found in Tables 7, 4, and 5.
3.3 Rule-based GEC for Zarma
As a baseline, we built a rule-based system using Levenshtein distance and a Bloom filter (Bloom, 1970), for Zarma (Figure 1). Additionally, we implemented a tool and API in Python. The tool—the first of its kind for Zarma, to our knowledge—is designed for a wide range of users. It offers command-line and graphical user interfaces (GUI) and is much less computationally intensive than the LLM-based approach. Moreover, our results show that it provides a competitive spell correction performance compared to the LLM- and MT-based approaches. To ensure the accessibility of the tool, we plan a public release upon acceptance.
Figure 1: Rule-Based GEC tool Workflow
The GEC process for Zarma text follows a structured sequence of steps, starting with the input text $T$, passing through a Bloom filter decision $B$ to determine if the word exists in the dictionary, leading to either a grammatical check or a correction algorithm, followed by a final dictionary lookup and correction $C$, as needed:
$C(T)=\begin{cases}T&\text{if }D(B(T))=\text{True},\\ \text{Correct}(T)&\text{if }D(B(T))=\text{False}.\end{cases}$ (1)
The detailed process includes:
Text Processing
The system begins by processing the text, where the Zarma content is segmented into words, punctuation, and spaces using regular expressions. This initial step ensures that every part of the text is ready for individual examination—a necessity given Zarma’s linguistic intricacies.
Bloom Filter
The Bloom filter—known for its space and time efficiency—performs probabilistic checks on whether a word might be in the dictionary. For a given word $w$, it uses a set of hash functions to probe various positions in a bit array. If all checked positions are flagged, $w$ might be in the dictionary:
$B(w)=\bigwedge_{i=1}^{k}\text{bit}[h_{i}(w)],$ (2)
If $B(T)=\text{False}$, the system assumes the word is not in the dictionary and immediately moves to the Correction Algorithm to suggest a correction. This process ensures efficiency when dealing with unknown words.
Dictionary Lookup
At the core of the system is the trie-based dictionary lookup. We built this dictionary using the Feriji dataset dictionary (Keita et al., 2024). If the Bloom Filter decision is positive ($B(T)=\text{True}$), the system proceeds to a Grammatical Check to verify if the word is grammatically correct. If the word is found to be grammatically incorrect, or if $B(T)=\text{False}$ (indicating the word is not in the dictionary), the system performs a Dictionary Lookup to confirm or correct the word.
Grammar Rules
The system incorporates a set of Zarma grammar rules, a task made challenging by the lack of a standard writing format. This phase scrutinizes elements like consonant rules and vowel lengths, essential for keeping the text accurate to Zarma’s grammatical essence despite its absence of a writing standard.
Correction Algorithm
When an error is detected, either because the Bloom Filter returned false or the grammatical check identified an issue, the Correction Algorithm is activated. This component leverages the Levenshtein distance $L$ to identify the smallest number of edits—insertions, deletions, or substitutions—required to rectify the erroneous word. For illustration, given a corrupted sentence "A sindq biri," the Levenshtein algorithm’s operation can be represented as:
Corrupted Sentence: "A sindq $\begin{array}[]{c}\text{{Corrupted Sentence}: "A sindq biri"}\\ \downarrow\\ \text{{LevSuggest}}("sindq")\\ \downarrow\\ \{\text{"sind"},\text{"sinda"}\}\end{array}$ (3)
Here, "sindq" undergoes comparison with the lexicon, and "sind" and "sinda" are suggested based on their proximity in terms of minimal edit distance. In this context, $L(a,b)$ calculates the distance between the incorrect string $a$ and a potential correct string $b$, considering their lengths $i$ and $j$, with $1_{(a_{i}\neq b_{j})}$ serving as an indicator function to highlight discrepancies. This mechanism ensures the algorithm’s efficiency in providing appropriate corrections, thus enhancing the GEC tool’s overall accuracy for Zarma text.
4 Data
We created a combined dataset of over 250,000 examples by merging two distinct datasets—a synthetic dataset and a human-annotated ”gold” dataset. We split it into 80% training, 10% validation, and 10% testing.
4.0.1 Synthetic Data
To address the scarcity of annotated data in Zarma, We applied a noise script to the Feriji dataset (Keita et al., 2024). This script introduces realistic typographical and grammatical errors into Zarma sentences, to mimick plausible human mistakes. The corruption process is guided by a defined set of operations to ensure the errors are representative of actual language use.
We applied a noise function N to each original sentence SZ (composed of a subject S, verb V, and complement C) to produce a corrupted sentence CSZ:
$CSZ(S,V,C)=N\circ SZ(S,V,C)$ (4)
The noise function N consists of four operations: deletions $\delta$, insertions $\mu$, substitutions $\sigma$, and transpositions $\tau$.
For example, consider the original sentence SZ = A go koy fuo” (He is going home”). Applying a substitution operation $\delta$ that changes ’go’ to ’ga’ yields the corrupted sentence CSZ = ’A ga koy fuo’. While ’ga’ is a valid Zarma word, in this context, it alters the meaning to ’He will go home,’ introducing a grammatical error.
$\textbf{SZ}="A\,go\,koy\,fuo"(I\ am\ going\ home)$ (5)
$\textbf{N}(\textbf{SZ})=\textbf{SZ}\xrightarrow{\sigma("go"\to"ga")}"A\,ga\,koy\,fuo"$
$\textbf{CSZ}="A\,ga\,koy\,fuo"(I\ will\ go\ home)$
[…]
Furthermore, to enrich the dataset with diverse linguistic nuances, the script generates four corrupted variants for each correct sentence in the dataset. A snapshot of the synthetic data is provided in Table 1.
Correct Sentence Corrupted Sentences
Sintina gaa Irikoy na beena da ganda taka. Sintina gaa Irikog na beena da ganda taka.
Sintina gaa Irikoy na been da ganda taka.
Sintina aga Irikoy na beena da ganda taka.
Sintina gaa Irikoy na beena ea ganda taka.
Table 1: Snapshot of The Synthetic Data
4.1 Human-Annotated Data
The second dataset was curated by human annotators proficient in Zarma. They introduced grammatical and logical errors (LG)—more details about this errors type can be found in Section B.2—into Zarma sentences and provided justifications for each alteration. This dataset exposes the models to a broader array of error types—particularly LG errors—and their corresponding corrections, enhancing the models’ capacity to generalize and accurately correct unseen texts in Zarma, especially in zero-shot or few-shot learning scenarios. An example from the human-annotated dataset is shown in Table 2
Incorrect Sentence Correction Error Explanation
Souba, Ay koy Niamey Souba, Ay ga koy Niamey "Souba" means tomorrow, and therefore the tense must be in the future using the future tense marker "ga" after the subject "Ay."
Table 2: Snapshot of the human-annotated dataset
5 Experiment
We conducted a series of experiments to evaluate the effectiveness of different GEC methods for Zarma. We selected three models for training based on their demonstrated proficiency in multilingual tasks and aptitude for few-shot learning: Gemma (Team et al., 2024), MT5-small, and M2M100. The training was conducted on Google Cloud, using a Deep Learning Virtual Machine equipped with NVIDIA GPUs T4 x2. Due to computational resource constraints, we applied QLoRA quantization (Dettmers et al., 2023) to the Gemma model to reduce memory requirements. For MT5-small and M2M100, we used smaller variants of the models to ensure feasible training times. For LLM-based methods (Gemma and MT5-small), we structured the combined dataset as described in Section 4, to incorporate both incorrect sentences and their corrections along with error explanations. For the MT-based method (M2M100), we adopted a MT task-specific format, using aligned pairs of incorrect and correct Zarma sentences without explicit error explanations. The detailed training settings for each model are presented in Table 7.
6 Results & Comparative Analysis
To assess the effectiveness of each GEC method for the types of errors addressed in this paper (as detailed in Section B), we conducted an evaluation consisting of two parts: an automatic evaluation (AE) and a manual evaluation (ME) by native Zarma speakers. We included the ME because AE tools are primarily developed and tested on high-resource languages. Therefore, their scores may not reliably reflect performance in low-resource settings such as Zarma. The ME allows us to obtain assurances from language speakers regarding the actual effectiveness of the methods.
6.1 Automatic Evaluation Results
In the AE, we employed two sets of metrics to evaluate our techniques. The first set consisted of common machine learning metrics, including the GLEU score (Wang et al., 2018) and the M2 score (Dahlmeier and Ng, 2012). However, these metrics are not applicable to the rule-based method due to its deterministic nature and the way these metrics are designed, which favor probabilistic models that produce varying outputs. The second set focused on spelling correction metrics—measuring the detection rate and suggestion accuracy. The results are presented in Table 3.
Methods AE Evaluation
Common ML Metrics Spelling Correction
GLEU Detection Correction
Rule-based - - 100% 96.27%
Gemma 2b 0.4572 0.528 76.19% 43.28%
MT5-small 0.6203 0.731 90.62% 57.15%
M2M100 0.7896 0.914 95.82% 78.90%
Table 3: AE metrics scores
6.1.1 Common Machine Learning Metrics
The GLEU and M2 scores provide an overall assessment of the models’ performance in correcting grammatical errors by comparing the corrected sentences to reference sentences. The higher scores achieved by the M2M100 model (GLEU: 0.7896, M2: 0.914) indicate its higher performance among the tested models. The LLM-based models—Gemma 2b and MT5-small—obtained moderate scores, which shows their ability to correct some errors but not as effectively as the M2M100 model.
6.1.2 Spelling Correction Metrics
For the spelling correction evaluation, we used a set of 3,539 sentences from the Feriji Dataset, introducing typographical errors using our corruption script (as described in Section 4). We measured two key metrics:
  • Detection Rate: The percentage of errors that the method successfully identified and attempted to correct.
  • Suggestion Accuracy: The percentage of corrections suggested by the method that were accurate.
As shown in Table 3, the rule-based method achieved a detection rate of 100% and a suggestion accuracy of 96.27%. This high performance is primarily due to its extensive dictionary and predefined rules, which allow it to effectively flag and correct errors covered within its scope. The M2M100 model also performed well, with a detection rate of 95.82% and a suggestion accuracy of 78.90%. This suggests that the MT-based approach is capable of effectively identifying and correcting spelling errors, even in a low-resource language setting. The LLM-based models—Gemma 2b and MT5-small—showed lower detection rates and suggestion accuracies. Gemma 2b detected 76.19% of errors with a suggestion accuracy of 43.28%, while MT5-small detected 90.62% of errors with a suggestion accuracy of 57.15%. These results indicate that while LLMs have some capacity to correct spelling errors, they are less effective than the MT-based approach and the rule-based method in this specific task.
6.2 Manual Evaluation Results
Given the limitations of AE metrics in low-resource settings, we conducted the ME involving five native Zarma speakers. The evaluators assessed how well each method corrected LG errors, using a scale from 1 to 5, where 1 indicates a poor correction and 5 indicates an excellent correction. The evaluation was performed on a set of 300 sentences from the Feriji Dataset, which were manually corrupted by our annotators to include LG errors. Detailed information about the evaluation protocol and the recruitment of evaluators can be found in Section C. The results are presented in Table 4.
Methods ME avg (1-5)
Logical Errors Correction Sentence Improvement
Rule-based 0.4 0
Gemma 2b 1 0
MT5-small 1.7 1
M2M100 3 2.5
Table 4: Manual Evaluation
The rule-based method struggled with LG errors, achieving low scores (0.4 for logical error correction and 0.0 for sentence improvement). This outcome is expected, as the rule-based approach relies on predefined rules and lacks the ability to understand the context or semantics of sentences. The M2M100 model outperformed the other methods, achieving higher scores in both logical error correction (3.0) and sentence improvement (2.5). This demonstrates that the MT-based approach is more effective at handling LG errors and improving the overall quality of the sentences. The LLM-based models achieved moderate scores, with MT5-small performing better than Gemma 2b. MT5-small received scores of 1.7 for logical error correction and 1.0 for sentence improvement, suggesting some capability in handling LG errors, but not as effectively as the M2M100 model. We observed that the models faced varying levels of difficulty with different types of LG errors. For instance, they were generally better at correcting verb tense errors than subject-verb agreement errors.
6.3 Zero-Shot scores
To further assess the generalization capabilities of our models, we conducted a zero-shot evaluation using a separate dataset. For that, We gathered a separate set of Zarma texts from books that had not been used during training. We then applied the same corruption script described in Section 4 and asked the annotators to include some LG errors. containing sentences with words and grammatical structures not present in the training data. This evaluation aimed to understand how well the models can handle new or unseen data, which is a common scenario in low-resource language settings. The results of the zero-shot evaluation are presented in Table 5.
Methods AE Evaluation Manual Evaluation avg (1-5)
Common ML Metrics Spelling Correction Logical Errors Correction Sentence Improvement
GLEU Detection Correction
Rule-based - - 94.17% 81.04% 0 0
Gemma 2b 0.3926 0.481 53.93% 26.52% 0.6 0
MT5-small 0.5018 0.657 73.64% 58.31% 0.8 1
M2M100 0.7683 0.904 92.27% 78.34% 2.7 1.3
Table 5: Performance Evaluation on Zero-Shot Dataset
As expected, the rule-based method showed a decrease in performance compared to the controlled evaluation. Its detection rate dropped to 94.17% and suggestion accuracy to 81.04%. This decline is due to the method’s reliance on a predefined dictionary and rules, which may not cover new words or grammatical constructions present in the zero-shot dataset. Additionally, the ME scored it with 0 points for both logical errors correction and sentence improvement.
The M2M100 model again demonstrated higher performance among the tested models in the zero-shot setting. It achieved a GLEU score of 0.7683 and an M² score of 0.904, indicating strong overall correction capability. Its detection rate was 92.27% with a suggestion accuracy of 78.34%, showing that it can effectively identify and correct errors in unseen data. Furthermore, in ME, it received 2.7 for logical errors correction and 1.3 for sentence improvement.
The LLM-based models, Gemma 2b and MT5-small, showed reduced performance in the zero-shot evaluation. Gemma 2b’s detection rate was 53.93% with a suggestion accuracy of 26.52%, while MT5-small achieved a detection rate of 73.64% and a suggestion accuracy of 58.31%. In terms of ME, Gemma 2b scored 0.6 for logical errors correction and 0 for sentence improvement, while MT5-small achieved 0.8 and 1, respectively.
Our results suggest that MT models trained on multilingual datasets hold significant potential for improving GEC in low-resource languages. The ME scores further validate that the M2M100 model excels not only in automated metrics but also in human judgments, particularly in complex correction tasks. To validate the reproducibility and robustness of our methods, we conducted additional experiments with Bambara—a language from a different linguistic family. The results of this experiment are detailed in Section A.
7 Potential Applications
Our team visited Niamey to present our work to the local Zarma community and gather their feedback. The discussions provided valuable insights into potential applications for our GEC tool and broader language models.
Content Creation and Educational Resources
One significant comment we received was the potential use of our model to translate coding content and general educational materials for enthusiasts and students. There is a growing interest in technology and programming within the community, but a considerable language barrier exists due to the lack of educational resources in Zarma. By translating and correcting coding tutorials, textbooks, and other educational materials into Zarma, our model can help overcome this challenge. Additionally, the GEC tool can be used to improve general educational content in Zarma, including textbooks, instructional materials, and online courses across various subjects.
Cultural Preservation and Documentation
Feedback from the community highlighted the importance of maintaining accurate written records of folklore, oral histories, and traditional knowledge. The GEC tool can support these efforts by providing a means for transcribing and publishing grammatically accurate texts.
8 Conclusion and Future Work
We presented a comparison of rule-based, LLM-based, and MT-based methods for Zarma GEC. The MT-based approach performed best in both controlled and zero-shot scenarios. We confirmed these findings by testing with Bambara language. This research establishes a foundation for future GEC development in other low-resource languages.
Future work includes hybrid models that combine rule-based precision with MT and LLM adaptability, data augmentation to improve performance, continuous learning mechanisms, resource optimization, and cross-lingual transfer learning to support a wider range of languages.
9 Limitations
9.1 Rule-Based Method
Rule-based approaches rely on predefined linguistic rules and dictionaries, which limit their adaptability. They do not handle new or unseen words well. In zero-shot scenarios, where the system encounters words or structures beyond its dictionary, it often fails to correct or detect errors. Additionally, the creation and maintenance of linguistic resources is difficult for low-resource languages, which often lack standardized orthography (Scannell, 2007).
9.2 LLMs Method
LLMs—Gemma 2b and MT5-small—must rely on diverse training data to capture linguistic nuances. Yet, existing LLMs are often trained on data from high-resource languages, which reduces performance in languages like Zarma. A lack of training samples in these languages causes the model to produce lower-quality suggestions (Bender et al., 2021). In addition, fine-tuning LLMs requires significant computational resources, which can be a challenge in low-resource settings.
9.3 MT Method
MT models—M2M100 in our studies—learn from parallel corpora that map incorrect text to correct text. In low-resource languages, such parallel data is often lacking or incomplete (Tiedemann, 2020). Even though multilingual models can leverage patterns from other languages, the absence of sufficient in-language data can lead to contextually imprecise corrections. Moreover, these models require targeted fine-tuning and periodic updates as language use evolves.
Appendix A Further Experiment with other Languages
After obtaining promising results for Zarma using LLM and MT-based approaches, we conducted further experiments to validate the reproducibility of these methods—using the M2M100 and MT5 models. We selected the Bambara language for this experiment because it belongs to a different linguistic family, which allows us to evaluate the performance of the approaches on a language outside the Nilo-Saharan family. We used the Bayelemabaga dataset (Vydrin et al., 2022) for this experiment. The same data preparation process described in the methodology section was followed; however, we excluded any human-annotated data to focus only on spelling correction performance. The results are presented in Table 6.
Methods AE Evaluation
Common ML Metrics Spelling Correction
GLEU Detection Correction
MT5-small 0.5713 0.643 87.45% 52.91%
M2M100 0.7293 0.851 94.64% 68.18%
Table 6: Performance Metrics for Bambara
The Bambara experiment demonstrated that the MT-based approach outperformed the LLMs-based one regarding word-level correction metrics. The MT-based approach achieved a detection rate of 94.64% and a suggestion accuracy of 68.18%. In contrast, the LLMs-based approach detected 87.45% of errors and suggested corrections with 52.91% accuracy. The results from the Bambara experiment highlight the potential of both LLMs and MT models to improve GEC for low-resource languages significantly. However, they also emphasize the necessity for continued expanding and diversifying training datasets.
Appendix B Errors Being Addressed
In this section, we explain the types of errors GEC methods address. We categorize the errors into two main types: word-level correction (spellchecking) and context-level correction. The context-level correction is further divided into grammar errors, logical errors, and sentence improvement. Below, we define each error type and provide examples to illustrate them.
B.1 Word-Level Correction (Spellchecking)
Word-Level correction involves identifying and correcting typographical errors in individual words. These errors are usually due to misspellings, incorrect usage of characters, or typographical mistakes.
  • Example:
    • Incorrect: Sintina gaa Irikog na beena da ganda taka.
    • Correct: Sintina gaa Irikoy na beena da ganda taka.
B.2 Context-Level Correction
Context-level correction involves errors that go beyond individual words and affect the overall structure and meaning of the sentence. We categorize these errors into logical errors and sentence improvement.
B.2.1 Logical Errors
Logical errors are inconsistencies in sentence meaning that arise from incorrect verb conjugations, subject-verb agreement, or the misuse of grammatical markers. Unlike grammatical errors, logical errors affect the semantic coherence of a sentence.
  • Example:
    • Incorrect: Souba, Ay koy Niamey. (The time indicator "Souba" means "tomorrow," but the verb "koy" indicates present tense.)
    • Correct: Souba, Ay ga koy Niamey. (The future tense marker "ga" matches the time indicator "Souba.")
B.2.2 Sentence Improvement
Sentence improvement focuses on enhancing readability, precision, and style. While the sentences may already be grammatically correct, improvements ensure better flow and clarity.
  • Example:
    • Original: I girbi honkuna i tun be.
    • Improved: I ga girbi suba.
Appendix C Evaluators Recruitment Process
To ensure a thorough and culturally appropriate manual evaluation of our GEC methods, we recruited five native Zarma speakers to serve as evaluators. The primary criteria for selection was their fluency in Zarma—their ability to read and write proficiently in the language. The evaluators were selected through community outreach in Niamey, Niger. All selected evaluators had completed secondary education or higher.
Prior to the evaluation, we conducted a training workshop to familiarize the evaluators with the objectives of the study, the evaluation protocols, and the scoring guidelines. During the evaluation process, each evaluator independently assessed a set of 300 sentences that had been manually corrupted to include logical and grammatical errors. They reviewed the corrections made by each GEC method and assigned scores based on accuracy and fluency, where:
  • 1 indicates a poor correction with significant errors remaining or introduced.
  • 5 indicates an excellent correction with all errors appropriately addressed and the sentence reading naturally.
All evaluators provided informed consent to participate in the study and were compensated for their time and contributions.
Appendix D Training Details
Models Parameters Training Details
QLoRA GPU Used Lr
Gemma 2b 2 billion Applied P100 $2\times 10^{-4}$
MT5-small 300 million Not Applied T4x2 $2\times 10^{-5}$
M2M100 418 million Not Applied P100 $2\times 10^{-5}$
Table 7: Training Settings for the models
Appendix E Other Figures
Figure 2: Images of the different GEC tool interfaces. The rule-based on the left and the other approaches on the right